# Roles of SN and AGN Feedback in Shaping the Baryonic Content in a Wide Range of Dark Matter Halo Mass

### v2

@@ Abstract
We build upon FEGA25 (Contini et al., 2025), a previously introduced semi-analytic model (SAM) for galaxy formation and evolution, by focusing on its enhanced treatment of supernova (SN) and active galactic nucleus (AGN) feedback mechanisms. In addition to the traditional AGN feedback mode—negative (suppressing cooling) and the new positive mode (triggering star formation)—we introduce two distinct implementations of a third mode: the ejection of hot gas beyond the virial radius (AGNeject1 and AGNeject2). This component addresses a longstanding issue in SAMs and hydrodynamical simulations: the overestimation of hot gas fractions in low- and intermediate-mass halos. FEGA25 is calibrated via MCMC using a suite of cosmological N-body simulations (YS50HR, YS200, YS300) and a comprehensive set of observed stellar mass functions across a wide redshift range. We find that SN feedback dominates gas ejection in halos with $\log M_{\rm halo}\lesssim 12$, while AGN feedback becomes increasingly important at higher halo masses. The AGNeject2 model, which activates primarily at late times ($z<1$), reproduces a characteristic “cavity” (sort of U-shape) in the baryon fraction at $z=0$, similar to trends observed in simulations like SIMBA and IllustrisTNG. Conversely, AGNeject1 yields a smoother, redshift-independent evolution. Both models preserve the stellar and cold gas components and successfully reproduce the stellar-to-halo mass relation up to $z=3$. Our results emphasize that a physically motivated AGN-driven mechanism capable of selectively removing hot gas is essential to model the baryon cycle accurately, particularly in the intermediate halo mass regime.
@@ 1 Introduction
In the current $\Lambda$CDM framework, galaxies form within dark matter (DM) halos that assemble hierarchically through mergers and smooth accretion (White & Rees, 1978; Springel et al., 2005). While the growth of DM structures is well described by gravitational dynamics, the observable properties of galaxies depend critically on baryonic processes—namely gas cooling, star formation, chemical enrichment, and feedback. Among these, feedback from supernovae (SN) and active galactic nuclei (AGN) is widely regarded as essential for regulating star formation, quenching massive galaxies, and explaining the observed galaxy stellar mass function (Guo et al., 2013; Somerville & Davé, 2015; Naab & Ostriker, 2017; Contini et al., 2019, 2020; Vogelsberger et al., 2020; Henriques et al., 2020).
@@
Yet, despite the inclusion of increasingly sophisticated feedback prescriptions, models still struggle to simultaneously match key observables such as the cosmic star formation rate density (Madau & Dickinson, 2014), the stellar-to-halo mass relation (Moster et al., 2013; Behroozi et al., 2019), and the distribution of hot and cold baryons across halo mass and redshift (Chiu et al., 2016; Eckert et al., 2021; Popesso et al., 2024). This tension is particularly severe in the group mass regime ($\log M_{\rm halo}\sim 12.5$–13.5), where feedback must be powerful enough to eject baryons, yet gentle enough not to alter the stellar and cold gas components.
@@
Large-scale hydrodynamic simulations such as IllustrisTNG (Pillepich et al., 2018; Nelson et al., 2019), SIMBA (Davé et al., 2019), EAGLE (Schaye et al., 2015), MAGNETICUM (Dolag et al., 2016), and more recently FLAMINGO (Schaye et al., 2023) and MillenniumTNG (Pakmor et al., 2023), have offered valuable insights into feedback-regulated galaxy evolution. However, they often yield divergent predictions for baryon fractions and hot gas content, reflecting the sensitivity of such quantities to the underlying physics. For instance, while IllustrisTNG and SIMBA predict a baryon “cavity”—a depression in the baryon fraction at intermediate halo masses—other simulations such as MAGNETICUM show a smoother, power-law trend more consistent with X-ray and SZ observations (Hadzhiyska et al., 2024; Popesso et al., 2024).
@@
Recent comparative analysis (Ayromlou et al., 2023; Salcido et al., 2023; Dev et al., 2024) have further demonstrated that baryon retention is highly dependent on the efficiency and coupling of feedback modes. These studies reinforce the need for physically motivated models that can selectively eject hot gas—especially in group-scale halos—without modifying the stellar and cold gas reservoirs (Eckert et al., 2021; Dev et al., 2024; Popesso et al., 2024).
@@
In this context, semi-analytic models (SAMs) remain an indispensable tool. Their modular structure and computational efficiency allow for rapid exploration of parameter space and clearer isolation of physical mechanisms. Modern SAMs such as those developed by Lagos et al. (2024) and Contini et al. (2025) have adopted increasingly sophisticated AGN feedback prescriptions, including both radio-mode heating and direct ejection of hot gas. Notably, the SHARK2 model (Lagos et al., 2024) includes a formulation similar in spirit to one of our schemes.
@@
Building on this foundation, we have developed the FEGA series (Contini et al., 2024, 2025), a SAM calibrated against a wide range of observational constraints. In particular, FEGA25 (Contini et al., 2025) introduced two key advancements: a physically motivated positive AGN feedback mode (triggering star formation under residual cooling), and a novel mechanism for ejecting hot gas beyond the virial radius. The latter is motivated by the need to reconcile models with observed hot gas fractions in low- and intermediate-mass halos, where supernova feedback alone appears insufficient (Henriques et al., 2013; Hirschmann et al., 2016; Ayromlou et al., 2023).
@@
In this work, we investigate the impact of two distinct implementations of this hot gas ejection mode—AGNeject1, which scales with black hole accretion and halo virial velocity, and AGNeject2, which uses the surplus AGN energy beyond that needed to suppress cooling to drive ejection. We calibrate both models using cosmological DM–only simulations and an extensive set of observed stellar mass functions from $z=3$ to the present (Marchesini et al., 2009, 2010; Ilbert et al., 2013; Muzzin et al., 2013; Tomczak et al., 2014; Bernardi et al., 2018).
@@
We focus on three main goals: (i) to quantify the relative contributions of SN and AGN feedback in ejecting baryons beyond the virial radius; (ii) to assess the evolution of baryon and hot gas fractions as a function of halo mass and redshift; and (iii) to compare our predictions with hydrodynamic simulations and observational estimates, particularly those from recent X-ray and SZ studies (Chiu et al., 2016; Eckert et al., 2021; Angelinelli et al., 2023; Popesso et al., 2024).
@@
Finally, in Section 4, we summarize our main findings and discuss their implications for galaxy formation theory and future observational efforts. Stellar masses are computed by adopting a Chabrier (2003) initial mass function, and unless otherwise specified, all masses are h-corrected.
@@ 2 Methods
To achieve our goals, we employ the state-of-the-art semi-analytic model FEGA25 (Formation and Evolution of GAlaxies Contini et al. 2025, hereafter C25), applied to a suite of three DM–only cosmological simulations. Additionally, we gain further insight from our high-resolution zoom-in simulation NewCluster (Han et al. 2025b). Below, we briefly outline the key features of our SAM, including the main physical prescriptions relevant to this study and, most notably, the new implementation of an alternative AGN-driven hot gas ejection mode, which represents a key innovation of this work. We also summarize the main characteristics of the NewCluster simulation. It is important to emphasize that in the analysis presented in Section 3, we make limited use of NewCluster predictions, as a dedicated and more comprehensive study based on this simulation is currently in preparation (Seo et al.).
@@ 2.1 FEGA25
FEGA25 is an updated version of the SAM originally introduced in Contini et al. 2024 (hereafter C24), featuring two major advancements: (i) a more sophisticated treatment of star formation that follows the Extended Kennicutt–Schmidt relation (Shi et al., 2011), and (ii) a novel implementation of positive AGN feedback, a mode that enhances star formation rather than suppressing it.
@@ 2.1.1 SN feedback
SN feedback, resulting from the explosive deaths of massive stars, is a fundamental process in galaxy formation and is central to the present model. This mechanism is particularly effective in regulating star formation in low-mass galaxies, shaping the low-mass end of the stellar mass function (SMF) and luminosity function (e.g., Guo et al. 2010; Henriques et al. 2013) by limiting stellar mass growth in dwarfs. The energy released—both radiative and kinetic—can substantially impact the interstellar medium, leading to its heating and ionization. In some cases, this energetic input is sufficient to drive gas completely out of the galactic halo, making SN feedback a critical component in virtually all SAMs of galaxy formation. Historically, SN feedback has been maximized in such models because it has been essential to reproducing the observed evolution of the SMF’s low-mass slope (e.g., Guo et al. 2011, 2013; Henriques et al. 2013, 2015; Hirschmann et al. 2016; Contini et al. 2017; Henriques et al. 2020).
In FEGA25, SN feedback is implemented such that cold gas in the galactic disk is heated and transferred to the hot gas phase, and part of the hot gas is further expelled into an external reservoir. This approach follows the formulation of Guo et al. (2011), though we adopt different parameters, following Hirschmann et al. (2016), in which the feedback weakens over cosmic time.
@@
where $\epsilon_{\rm{reh}}=0.7$, $V_{\rm{reh}}=70\,\rm{km/s}$, $\beta_{\rm{reh}}=3.5$, and $\delta M_{*}$ represents the mass of newly formed stars. The inclusion of the variable $\epsilon_{\rm{reh}}$ by Guo et al. (2011) allows for enhanced ejection efficiencies in dwarf galaxies. The total energy injected into the disk and halo gas is:
@@
where $\eta_{\rm{ej}}=0.15$, $V_{\rm{ej}}=70\,\rm{km/s}$, $\beta_{\rm{ej}}=3.5$, and $0.5V_{\rm{SN}}^{2}$ represents the average kinetic energy of SN ejecta per unit mass of stars formed, calculated based on $V_{\rm{SN}}=630$ km/s (Croton et al. 2006). From this energy amount, the ejected gas (including metals) from the halo is:
@@ 2.1.2 Negative and Positive AGN feedback
The radio mode of AGN feedback in our model follows the original prescription of Croton et al. (2006). In this mode, hot gas accretes onto the central black hole (BH), releasing mechanical energy that heats the surrounding halo gas. The accretion rate is given by:
@@
Introduced in C24, the positive AGN feedback simulates a secondary starburst triggered by AGN-driven pressure. This effect, while supported by both observational (e.g., Cresci et al. 2015; Mahoro et al. 2017; Joseph et al. 2022) and theoretical studies (Gaibler et al. 2012; Silk 2013; Mukherjee et al. 2018), was initially modeled in a simplified way to test its impact, but introducing new free parameters.
@@ 2.1.3 AGNeject1 and AGNeject2
where $V_{\rm{scale}}$ is a tunable parameter determined during model calibration, and it sets the threshold at which this ejection becomes efficient  [^1] This prescription remains valid even when $\dot{M}_{\rm{cool}}>2\frac{\dot{E}{\rm{radio}}}{V_{200}^{2}}$, i.e., when the AGN energy input is insufficient to fully suppress gas cooling.. High values of $V_{\rm{scale}}$ make gas ejection more effective in massive halos, whereas low values diminish efficiency in low-mass halos. The computed ejected mass, $M_{\rm{ejected}}$, is stored in an external reservoir but remains eligible for future reincorporation, consistent with the treatment described in C24. Naturally, when $V_{200}>V_{\rm{scale}}$, the ejected mass is set to zero.
@@
Analogously to the implementation in AGNeject1, we assume that gas ejection becomes progressively more challenging in more massive halos—specifically, it is governed by the virial velocity of the host halo. Consequently, in cases where $\dot{M}_{\rm{cool}}<2\frac{\dot{E}{\rm{radio}}}{V_{200}^{2}}$, the surplus energy is assumed to expel a portion of the excess gas mass, $M_{\rm{excess}}$, which would otherwise have been reheated to the virial temperature. The amount of ejected mass is computed using the following expression:
@@
In summary, the full AGN feedback framework implemented in FEGA25 consists of three components: the negative mode, the positive mode, and the hot gas ejection mode, either AGNeject1 or AGNeject2. Compared to the version presented in C24, FEGA25 introduces a more physically motivated description of the positive feedback (see Equation 7) and adds a new prescription for the ejection of hot gas (Equations 8 or  9). As a result, FEGA25 stands out as the only SAM that incorporates both a treatment of AGN-driven positive feedback and a mechanism for the ejection of hot gas. It must be noted, however, that an implementation similar to AGNeject2 is incorporated in SHARK2 SAM (Lagos et al., 2024).
@@ 2.1.4 Set of Simulations and Calibration
Calibrating a SAM is a fundamental step, particularly given that the physical prescriptions governing galaxy formation typically involve numerous parameters that are either loosely constrained by observations or entirely unknown. We calibrate FEGA25 following the same procedure adopted in C25—namely, employing a Markov Chain Monte Carlo (MCMC) method. For a comprehensive explanation of this method, including its strengths, limitations, and a discussion of alternative strategies, we refer the reader to C24, Henriques et al. (2020), and references therein.
The calibration of FEGA25 relies on the same compilation of observed SMFs used in C24 and C25, spanning redshifts from $z=3$ down to $z=0$. This dataset brings together measurements from several observational studies, offering a robust picture of the SMF over time. Specifically, we use data points at $z=3$, $z=2.75$, $z=2$, $z=1.75$, $z=1$, $z=0.75$, $z=0.4$, and $z=0$, with high-redshift SMFs taken from Marchesini et al. (2009), Marchesini et al. (2010), Ilbert et al. (2010), Sánchez et al. (2012), Muzzin et al. (2013), Ilbert et al. (2013), and Tomczak et al. (2014). The SMF at $z=0$ is compiled from Baldry et al. (2008), Li & White (2009), Baldry et al. (2012), and Bernardi et al. (2018). To complete the calibration pipeline and generate galaxy catalogs, we make use of cosmological N-body simulations, which provide the merger trees—the structural backbone of the SAM—containing essential information about the evolution of DM halos where galaxies reside.
We utilize three simulations: the YS50HR box, with a comoving volume of $(50\,\rm{Mpc/h})^{3}$, is used during the initial calibration phase, a larger YS200 box, with a volume of $(200\,\rm{Mpc/h})^{3}$, is employed for the final calibration, and a new run YS300, with a volume of $(300\,\rm{Mpc/h})^{3}$. All simulations cover the redshift range from $z=63$ to $z=0$, with 100 snapshots distributed between $z=20$ and the present. They are run using GADGET4 (Springel et al. 2021), adopting cosmological parameters from the Planck 2018 results (Planck Collaboration et al. 2020): $\Omega_{m}=0.31$ (total matter density), $\Omega_{\Lambda}=0.69$ (dark energy density), $n_{s}=0.97$ (scalar spectral index), $\sigma_{8}=0.81$ (amplitude of matter fluctuations), and $h=0.68$ (Hubble parameter normalized to $100\,\rm{km\,s^{-1}\,Mpc^{-1}}$). The corresponding mass resolutions are $M_{\rm{DM,part}}=10^{7}\,M_{\odot}/h$, $3.26\cdot 10^{8}\,M_{\odot}/h$, and $2.2\cdot 10^{9}\,M_{\odot}/h$ for the YS50HR, YS200, and YS300 simulations, respectively. For additional details on the overall setup and construction of the simulations, we refer the reader to Contini et al. (2023).
In line with the approach adopted by C25, the parameters constrained during the calibration process are: the AGN efficiency parameter $\kappa_{\rm{AGN}}$ appearing in Equation 4, the scale velocity $V_{\rm{scale}}$ in Equation 8, and the reincorporation factor $\gamma_{2}$. [^2] This parameter governs the rate at which ejected gas returns to the hot gas phase within the virial radius, and is defined by the relation: $\dot{M}_{\rm{ej}}=\gamma_{2}\frac{M_{\rm{ej}}}{t_{\rm{reinc}}}$, where $M_{\rm{ej}}$ represents the amount of gas in the ejected reservoir, and $t_{\rm{reinc}}$ is the reincorporation timescale. Further details are available in C24. The calibration of the SAM using the AGNeject1 prescription has already been performed in C25; therefore, in this work, we focus solely on calibrating the model with the AGNeject2 option enabled. The calibrated parameter values obtained are as follows for AGNeject1 (AGNeject2): $\kappa_{\rm{AGN}}=2\times 10^{-5},(5\times 10^{-5})$ $M_{\odot}$/yr; $V_{\rm{scale}}=500,(415)$ km/s; and $\gamma_{2}=0.378,(0.367)$.
We run the models on the merger trees extracted from all three simulations to generate our final galaxy catalogs. Due to the varying resolutions of the simulations, we apply different mass thresholds when selecting halos: specifically, we consider halos with $\log M_{200}>11.0$ in YS50HR, $\log M_{200}>12.0$ in YS200, and $\log M_{200}>14.0$ in YS300. The combination of these subcatalogs forms our final, comprehensive catalogs.

### v4

# Roles of Supernova and Active Galactic Nucleus Feedback in Shaping the Baryonic Content in a Wide Range of Dark Matter Halo Masses

@@ Abstract
We build upon FEGA25 (E. Contini et al., 2025), a previously introduced semi-analytic model (SAM) for galaxy formation and evolution, by focusing on its enhanced treatment of supernova (SN) and active galactic nucleus (AGN) feedback mechanisms. In addition to the traditional AGN feedback mode—negative (suppressing cooling) and the new positive mode (triggering star formation)—we introduce two distinct implementations of a third mode: the ejection of hot gas beyond the virial radius (AGNeject1 and AGNeject2). This component addresses a longstanding issue in SAMs and hydrodynamical simulations: the overestimation of hot gas fractions in low- and intermediate-mass halos. FEGA25 is calibrated via MCMC using a suite of cosmological N-body simulations (YS50HR, YS200, YS300) and a comprehensive set of observed stellar mass functions across a wide redshift range. We find that SN feedback dominates gas ejection in halos with $\log M_{\rm halo}\lesssim 12$, while AGN feedback becomes increasingly important at higher halo masses. The AGNeject2 model, which activates primarily at late times ($z<1$), reproduces a characteristic “cavity” (sort of U-shape) in the baryon fraction at $z=0$, similar to trends observed in simulations like SIMBA and IllustrisTNG. Conversely, AGNeject1 yields a smoother, redshift-independent evolution. Both models preserve the stellar and cold gas components and successfully reproduce the stellar-to-halo mass relation up to $z=3$. Our results emphasize that a physically motivated AGN-driven mechanism capable of selectively removing hot gas is essential to model the baryon cycle accurately, particularly in the intermediate halo mass regime.
@@
show]emanuele.contini82@google.com
show]yi@yonsei.ac.kr
@@ 1 Introduction
[…]
@@
Yet, despite the inclusion of increasingly sophisticated feedback prescriptions, models still struggle to simultaneously match key observables such as the cosmic star formation rate density (P. Madau & M. Dickinson, 2014), the stellar-to-halo mass relation (B. P. Moster et al., 2013; P. Behroozi et al., 2019), and the distribution of hot and cold baryons across halo mass and redshift (I. Chiu et al., 2016; D. Eckert et al., 2021; P. Popesso et al., 2024). This tension is particularly severe in the group mass regime ($\log M_{\rm halo}\sim 12.5$–13.5), where feedback must be powerful enough to eject baryons, yet gentle enough not to alter the stellar and cold gas components.
@@
Large-scale hydrodynamic simulations such as IllustrisTNG (A. Pillepich et al., 2018; D. Nelson et al., 2019), SIMBA (R. Davé et al., 2019), EAGLE (J. Schaye et al., 2015), MAGNETICUM (K. Dolag et al., 2016), and more recently FLAMINGO (J. Schaye et al., 2023) and MillenniumTNG (R. Pakmor et al., 2023), have offered valuable insights into feedback-regulated galaxy evolution. However, they often yield divergent predictions for baryon fractions and hot gas content, reflecting the sensitivity of such quantities to the underlying physics. For instance, while IllustrisTNG and SIMBA predict a baryon “cavity”—a depression in the baryon fraction at intermediate halo masses—other simulations such as MAGNETICUM show a smoother, power-law trend more consistent with X-ray and SZ observations (B. Hadzhiyska et al., 2024; P. Popesso et al., 2024).
@@
Recent comparative analysis (M. Ayromlou et al., 2023; J. Salcido et al., 2023; A. Dev et al., 2024) have further demonstrated that baryon retention is highly dependent on the efficiency and coupling of feedback modes. These studies reinforce the need for physically motivated models that can selectively eject hot gas—especially in group-scale halos—without modifying the stellar and cold gas reservoirs (D. Eckert et al., 2021; A. Dev et al., 2024; P. Popesso et al., 2024).
@@
In this context, semi-analytic models (SAMs) remain an indispensable tool. Their modular structure and computational efficiency allow for rapid exploration of parameter space and clearer isolation of physical mechanisms. Modern SAMs such as those developed by C. d. P. Lagos et al. (2024) and E. Contini et al. (2025) have adopted increasingly sophisticated AGN feedback prescriptions, including both radio-mode heating and direct ejection of hot gas. Notably, the SHARK2 model (C. d. P. Lagos et al., 2024) includes a formulation similar in spirit to one of our schemes.
@@
Building on this foundation, we have developed the FEGA series (E. Contini et al., 2024, 2025), a SAM calibrated against a wide range of observational constraints. In particular, FEGA25 (E. Contini et al., 2025) introduced two key advancements: a physically motivated positive AGN feedback mode (triggering star formation under residual cooling), and a novel mechanism for ejecting hot gas beyond the virial radius. The latter is motivated by the need to reconcile models with observed hot gas fractions in low- and intermediate-mass halos, where supernova feedback alone appears insufficient (B. M. B. Henriques et al., 2013; M. Hirschmann et al., 2016; M. Ayromlou et al., 2023).
@@
In this work, we investigate the impact of two distinct implementations of this hot gas ejection mode—AGNeject1, which scales with black hole accretion and halo virial velocity, and AGNeject2, which uses the surplus AGN energy beyond that needed to suppress cooling to drive ejection. We calibrate both models using cosmological DM–only simulations and an extensive set of observed stellar mass functions from $z=3$ to the present (D. Marchesini et al., 2009, 2010; O. Ilbert et al., 2013; A. Muzzin et al., 2013; A. R. Tomczak et al., 2014; M. Bernardi et al., 2018).
@@
We focus on three main goals: (i) to quantify the relative contributions of SN and AGN feedback in ejecting baryons beyond the virial radius; (ii) to assess the evolution of baryon and hot gas fractions as a function of halo mass and redshift; and (iii) to compare our predictions with hydrodynamic simulations and observational estimates, particularly those from recent X-ray and SZ studies (I. Chiu et al., 2016; D. Eckert et al., 2021; M. Angelinelli et al., 2023; P. Popesso et al., 2024).
@@
Finally, in Section 4, we summarize our main findings and discuss their implications for galaxy formation theory and future observational efforts. Stellar masses are computed by adopting a G. Chabrier (2003) initial mass function, and unless otherwise specified, all masses are h-corrected.
@@ 2 Methods
To achieve our goals, we employ the state-of-the-art semi-analytic model FEGA25 (Formation and Evolution of GAlaxies E. Contini et al. 2025, hereafter C25), applied to a suite of three DM–only cosmological simulations. Additionally, we gain further insight from our high-resolution zoom-in simulation NewCluster (S. Han et al. 2025b). Below, we briefly outline the key features of our SAM, including the main physical prescriptions relevant to this study and, most notably, the new implementation of an alternative AGN-driven hot gas ejection mode, which represents a key innovation of this work. We also summarize the main characteristics of the NewCluster simulation. It is important to emphasize that in the analysis presented in Section 3, we make limited use of NewCluster predictions, as a dedicated and more comprehensive study based on this simulation is currently in preparation (Seo C., et al.).
@@ 2.1 FEGA25
FEGA25 is an updated version of the SAM originally introduced in E. Contini et al. 2024 (hereafter C24), featuring two major advancements: (i) a more sophisticated treatment of star formation that follows the Extended Kennicutt–Schmidt relation (Y. Shi et al., 2011), and (ii) a novel implementation of positive AGN feedback, a mode that enhances star formation rather than suppressing it.
@@ 2.1.1 SN feedback
SN feedback, resulting from the explosive deaths of massive stars, is a fundamental process in galaxy formation and is central to the present model. This mechanism is particularly effective in regulating star formation in low-mass galaxies, shaping the low-mass end of the stellar mass function (SMF) and luminosity function (e.g., Q. Guo et al. 2010; B. M. B. Henriques et al. 2013) by limiting stellar mass growth in dwarfs. The energy released—both radiative and kinetic—can substantially impact the interstellar medium, leading to its heating and ionization. In some cases, this energetic input is sufficient to drive gas completely out of the galactic halo, making SN feedback a critical component in virtually all SAMs of galaxy formation. Historically, SN feedback has been maximized in such models because it has been essential to reproducing the observed evolution of the SMF’s low-mass slope (e.g., Q. Guo et al. 2011, 2013; B. M. B. Henriques et al. 2013, 2015; M. Hirschmann et al. 2016; E. Contini et al. 2017; B. M. B. Henriques et al. 2020).
In FEGA25, SN feedback is implemented such that cold gas in the galactic disk is heated and transferred to the hot gas phase, and part of the hot gas is further expelled into an external reservoir. This approach follows the formulation of Q. Guo et al. (2011), though we adopt different parameters, following M. Hirschmann et al. (2016), in which the feedback weakens over cosmic time.
@@
where $\epsilon_{\rm{reh}}=0.7$, $V_{\rm{reh}}=70\,\rm{km/s}$, $\beta_{\rm{reh}}=3.5$, and $\delta M_{*}$ represents the mass of newly formed stars. The inclusion of the variable $\epsilon_{\rm{reh}}$ by Q. Guo et al. (2011) allows for enhanced ejection efficiencies in dwarf galaxies. The total energy injected into the disk and halo gas is:
@@
where $\eta_{\rm{ej}}=0.15$, $V_{\rm{ej}}=70\,\rm{km/s}$, $\beta_{\rm{ej}}=3.5$, and $0.5V_{\rm{SN}}^{2}$ represents the average kinetic energy of SN ejecta per unit mass of stars formed, calculated based on $V_{\rm{SN}}=630$ km/s (D. J. Croton et al. 2006). From this energy amount, the ejected gas (including metals) from the halo is:
@@ 2.1.2 Negative and Positive AGN feedback
The radio mode of AGN feedback in our model follows the original prescription of D. J. Croton et al. (2006). In this mode, hot gas accretes onto the central black hole (BH), releasing mechanical energy that heats the surrounding halo gas. The accretion rate is given by:
@@
Introduced in C24, the positive AGN feedback simulates a secondary starburst triggered by AGN-driven pressure. This effect, while supported by both observational (e.g., G. Cresci et al. 2015; A. Mahoro et al. 2017; P. Joseph et al. 2022) and theoretical studies (V. Gaibler et al. 2012; J. Silk 2013; D. Mukherjee et al. 2018), was initially modeled in a simplified way to test its impact, but introducing new free parameters.
@@ 2.1.3 AGNeject1 and AGNeject2
where $V_{\rm{scale}}$ is a tunable parameter determined during model calibration, and it sets the threshold at which this ejection becomes efficient  [^1] This prescription remains valid even when $\dot{M}_{\rm{cool}}>2\frac{\dot{E}_{\rm{radio}}}{V_{200}^{2}}$, i.e., when the AGN energy input is insufficient to fully suppress gas cooling.. High values of $V_{\rm{scale}}$ make gas ejection more effective in massive halos, whereas low values diminish efficiency in low-mass halos. The computed ejected mass, $M_{\rm{ejected}}$, is stored in an external reservoir but remains eligible for future reincorporation, consistent with the treatment described in C24. Naturally, when $V_{200}>V_{\rm{scale}}$, the ejected mass is set to zero.
@@
Analogously to the implementation in AGNeject1, we assume that gas ejection becomes progressively more challenging in more massive halos—specifically, it is governed by the virial velocity of the host halo. Consequently, in cases where $\dot{M}_{\rm{cool}}<2\frac{\dot{E}_{\rm{radio}}}{V_{200}^{2}}$, the surplus energy is assumed to expel a portion of the excess gas mass,
$M_{\rm{excess}}=\left(2\frac{\dot{E}_{\rm{radio}}}{V_{200}^{2}}-\dot{M}_{\rm{cool}}\right)\cdot\rm dt,$
which would otherwise have been reheated to the virial temperature. The amount of ejected mass is computed using the following expression:
@@
In summary, the full AGN feedback framework implemented in FEGA25 consists of three components: the negative mode, the positive mode, and the hot gas ejection mode, either AGNeject1 or AGNeject2. Compared to the version presented in C24, FEGA25 introduces a more physically motivated description of the positive feedback (see Equation 7) and adds a new prescription for the ejection of hot gas (Equations 8 or  9). As a result, FEGA25 stands out as the only SAM that incorporates both a treatment of AGN-driven positive feedback and a mechanism for the ejection of hot gas. It must be noted, however, that an implementation similar to AGNeject2 is incorporated in SHARK2 SAM (C. d. P. Lagos et al., 2024).
@@ 2.1.4 Set of Simulations and Calibration
Calibrating a SAM is a crucial step, as the physical prescriptions governing galaxy formation involve many parameters that are often poorly constrained. We calibrate FEGA25 using the same MCMC approach as in C25. For a broader discussion of this method, including its advantages and limitations, we refer to C24, B. M. B. Henriques et al. (2020), and references therein.
The calibration relies on the same compilation of observed SMFs used in C24 and C25, spanning $z=3$ to $z=0$, with data from D. Marchesini et al. (2009, 2010); O. Ilbert et al. (2010); S. F. Sánchez et al. (2012); A. Muzzin et al. (2013); O. Ilbert et al. (2013); A. R. Tomczak et al. (2014) at high redshift, and I. K. Baldry et al. (2008); C. Li & S. D. M. White (2009); I. K. Baldry et al. (2012); M. Bernardi et al. (2018) at $z=0$. These datasets provide a robust description of the SMF evolution. To build galaxy catalogs, we use cosmological N-body simulations that supply the merger trees hosting galaxies.
Three simulations are employed: YS50HR ($50\,\rm{Mpc/h}$ box), YS200 ($200\,\rm{Mpc/h}$), and YS300 ($300\,\rm{Mpc/h}$). All run with GADGET4 (V. Springel et al., 2021) from $z=63$ to $z=0$ (100 snapshots between $z=20$ and 0), adopting Planck 2018 cosmology (Planck Collaboration et al., 2020). Their DM particle masses are $10^{7}$, $3.26\times 10^{8}$, and $2.2\times 10^{9}\,M_{\odot}/h$, respectively (see E. Contini et al. 2023 for details).
[…]
Finally, the models are run on merger trees from all three simulations to construct the catalogs. Due to different resolutions, we select halos above $\log M_{200}=11.0$ (YS50HR), 12.0 (YS200), and 14.0 (YS300). These combined subcatalogs form our final dataset for analysis.